An oblong whose dimensions are an exact number of cm has the same area in cm2 as its perimeter in cm. What are its dimensions?
1
Expert's answer
2011-06-14T10:01:13-0400
Its dimensions are a=3 cm and b=6 cm.
Perimeter is 2a+2b=6+12=18cm, area is 3*6=18 cm2.